OpenStudy (anonymous):

simplify: 5t-10 / t^2 -4

OpenStudy (aroub):

=5(t-2)/ (t+2)(t-2) =5/t+2

OpenStudy (anonymous):

\[\frac{5t-10}{t^2-4}=\frac{5(t-2)}{(t-2)(t+2)}=\frac{5}{t+2}\]

OpenStudy (bme89):

5/(t+2) where t may not equal 2

OpenStudy (aroub):

yes and u have to write that the must not equal to -2

OpenStudy (aroub):

t is not equal to -2 ***

OpenStudy (anonymous):

\[t \neq -2\]

OpenStudy (aroub):

exactly! :)

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

ty, aroub give one take one ,here you go

OpenStudy (aroub):

hehehe thank you so much!!! :D :D

OpenStudy (anonymous):

hehehehehehehehehe